# Neutron Reflectometry by Gradient Descent

### v1

@@ 1.5 Related work
In several works, neural-network surrogates have been proposed as an alternative to optimisation for the NR inverse problem in order to extract unknown parameters directly from data. Methods have been proposed in both NR [19] and X-ray diffraction [21], with the latter motivated as a way to provide real-time process control based on surrogate inversions. A recent review of ML technology applied to scattering problems can be found in [20]. In such cases the principal idea is to pre-train a neural-network surrogate on indicative reflectometry curves (either simulated or experimental) such that it is able to directly extract the parameters of interest. The key advantage is that for a new unseen testing reflectometry curve, the parameters can be extracted in ‘one-shot’ without the need for potentially computationally expensive algorithms when solving the NR inverse problem. While such approaches have merit, it is the opinion of the authors that they are weakened by the requirement to train the methods on large quantities of appropriate labelled training data (experimental or otherwise). While there are certainly situations in which this will be plausible it represents a major limitation to the practicality of such methods.
Recently, surrogate modelling has been proposed in an inference setting for NR. In [22], Starostin et al. propose to apply techniques from neural posterior estimation and prior amortisation to train a neural network to produce credible prior distributions based on reflectivity curves. The idea is that the network is trained to rapidly identify credible regions of the parameter space in the form of prior distributions that can be supplied to traditional Bayesian inference techniques such as Markov-chain Monte-Carlo (MCMC) or importance sampling (IS). The result is a significant reduction in the cost of evaluating a posterior over the NR parameters. The results of the paper are impressive and show that credible prior distributions can be found that include the multimodality that occurs due to the phase invariance problem when using very wide priors. As with all simulation based modelling approaches however, the requirement to pre-train a model (in this case on simulated data from a fixed number of layers) remains a limitation of the approach.
@@
Rather than implement any single bespoke gradient-based optimisation or inference technique for NR, the authors present here an open-source software library for fast computation of forward reflectivity and gradients in the python language. The library, titled refjax, is available at github.com/MDCHAMP/refjax. The library is built on top of the python JAX ecosystem from Google [34] that provides flexible automatic differentiation, parallelism and GPU acceleration. refjax provides simple access to the forward specular reflectivity calculation (including smearing) for an arbitrary slab model. In refjax, all input parameters are collected in standard python dictionaries with a standardised format.
@@
A key advantage of the provided library is the freedom to apply any optimisation routine developed in the JAX ecosystem and beyond. To date, highly performant packages are available for traditional optimisers (LM, BFGS) [35], (stochastic) gradient descent [36], HMC (including the no U-turn sampler [29]) [37] and others. Several examples of how to use the library in conjunction with popular ML optimisation suites are provided at https://github.com/MDCHAMP/refjax/tree/main/examples.
@@ 3.1 Crystalline quartz
In the first case study, the performance of gradient based techniques are highlighted on an existing dataset. To this end, the parameters of a highly homogeneous crystal quartz film deposited on a flat silicon wafer with a thin native oxide layer are considered. The reflectivity data were collected on the D17 neutron reflectivity instrument at the ILL (Grenoble, France) [38]. As part of the data collection, pointwise full-width at half maximum (FWHM) smearing kernel estimates were estimated. These were used to perform smearing in the forward reflectivity model as fixed known parameters. The slab model considered in this case study is provided in Table 1 and the NR parameters are displayed in Table 2. All parameters are considered unknown and are fit using both the optimisation and inference methods.
@@
First, an optimisation approach is considered. To perform the model fit, the widely-used stochastic gradient descent technique ADAM [24] is employed. Starting from an approximate initial values for the relevant parameters (listed in Tables 1 and 2), the optimiser descends the gradient of a $\chi^{2}$ error metric,
@@
for 2000 iterations with a learning rate of 0.05. The resulting fit to the data is given in Figure 3. As can be seen in the Figure, the fit to the data is excellent, achieving a $\chi^{2}$ error of 1.302. This is marginally lower than the value of $\chi^{2}=1.32$ reported in [38], indicating that the approach has found optimal parameters. Note that the best fit parameters for the native silicon thickness, SLD and the substrate SLD do not correspond to physical values in this case, as the adjacent roughnesses are very large. These are just fitting parameters to produce the final best fit SLD profile using a slab model.
Figure 3: Optimisation of quartz thin film NR dataset fit using gradient decent with the ADAM optimiser [24]. All parameters are floating and so free to fit. The identified $\chi^{2}$ error is lower than the value of $\chi^{2}=1.32$ reported previously in [38] for the same data.
With the effectiveness of gradient based optimisation established, a further study considering gradient-powered inference for NR fitting is presented here. The same quartz data is used to perform Bayesian inference of the unknown parameters by HMC. The NUTS sampler [29] (as implemented in the ‘numpyro’ library [37]) is used to sample from the posterior distribution over the parameters. NUTS has the advantage that the hyperparameter $L$ is set automatically in each iteration leading to efficient exploration of the typical set. The other HMC parameters are adapted using the default strategy in the numypro library for a burn-in period of 1000 samples.
The prior distributions were set to be uniform in the vicinity of the initial parameters used by the gradient scheme (see Tables 1 and 2). The chain was initialised prior to the burn-in at the optimal value found by gradient descent. Overall, 10 independent chains were run, each comprising 2000 samples drawn after burn-in. Posterior means and standard deviations are recorded in Tables 1 and 2. The resulting predictive posterior $p(\hat{R}|R)$ is plotted in Figure 4, wherein an excellent fit to the data is seen. The mean prediction of the posterior $E[\hat{R}$] has a $\chi^{2}$ error of 1.22, an improvement over previously published results ($\chi^{2}=1.32$, [38]) although we note that in our work, all parameters are left free to fit. In Figure 5, the marginal posterior distributions of the NR parameters of interest are shown alongside the distribution over the SLD profile of the quartz film.
@@
It is noteworthy that the obtained results are obtained by drawing only 2000 samples from the NUTS algorithm per chain, a fraction of the number of samples typically required by MCMC methods for good convergence. This phenomenon is in line with the efficiency gains expected from HMC over traditional MCMC methods, as demonstrated in other fields [39, 40]. However, it is of interest to quantify this phenomenon in the context of NR inversion.
In order to compare the efficiency of HMC (here NUTS) against gradient-free MCMC methods, Figure 6 plots the effective sample size (ESS) against iteration for both HMC and a Sample-adaptive MCMC (SA-MCMC) approach [41] implemented in the numpyro library [37]. ESS gives a measure of how many independent samples a correlated Markov-chain is equivalent to. A common rule of thumb in the Bayesian inference literature is that an ESS value of 200 per parameter represents a good coverage of the posterior for well-converged chains, whereas low values can correspond to poor exploration and autocorrelation.
@@
Figure 6: Comparison of effective sample sizes per iteration between a Hamiltonian Monte-Carlo (NUTS) scheme (green) and a gradient-free sample adaptive MCMC approach (orange). Each line is the effective sample size of a single parameter in a single chain. The horizontal line corresponds to a sample size of 200 which is a common rule of thumb for well converged chains in the literature [42].
As can be seen in Figure 6, the NUTS samples increase in ESS broadly monotonically with iteration, indicating good convergence and efficient exploration of the typical set even after only 1000 samples. SA-MCMC appears to start well (by exploring a local region of the typical set efficiently) but slow mixing and high auto-correlation lead to a decrease in ESS. It is interesting to note that although the per-sample computational complexity for NUTS is several times higher than SA-MCMC, better exploration of the typical set is seen with the gradient-based scheme, in line with results from other fields [39, 40].
@@ 3.2 Organic LED
To produce an approximate inference over the co-fit posterior of all four devices a VI scheme is adopted. The surrogate posterior $q$ is selected to be a multivariate normal distribution defined in an unconstrained space. That is, all parameters are mapped from their physical bounded space to one that is unconstrained by an appropriate invertible transformation. The resulting ELBO is evaluated by automatic differentiation (using the numpyro library) and the refjax software for reflectivity kernel calculations. The mean of the surrogate posterior is initialised at identical values for each device and the variance is initialised as a diagonal matrix $\sigma_{0}I$ with $\sigma_{0}=0.1$. The ELBO is optimised using the ADAM [24] optimiser for 2000 iterations with a learning rate of 0.05. To ensure that the optimisation does not become trapped in local minima, the optimisation procedure is run 100 times and the best (largest) value of the ELBO is retained as surrogate posterior. VI has significant computational advantages over sampling based approaches, each run of the inference is complete in <20 seconds on our hardware (AMD Ryzen Threadripper PRO 5995WX). Successive restarts can be run entirely in parallel and so the VI approach is very attractive with regard to modern CPU and GPU acceleration.
@@ 4 Discussion
In the second case-study, techniques from variational inference show remarkable performance in a complex NR fitting task. The authors believe that VI can be revolutionary for several NR applications particularly in settings whereby uncertainty quantification of high-throughput data is critical. An example is fast kinetic NR experiments whereby large quantities of NR data are collected dynamically and in-situ inference of model parameters is required. A key application of VI in NR might be in real-time parameter inference for closed-loop process control, something that has thus far required fast model surrogates and has been unattainable with traditional inverse modelling techniques. Another application could be in iterative experimental design whereby the parameter posterior distributions of initial experiments are used to inform the design of subsequent experiments during a single beamtime. Here, techniques from Bayesian optimisation [43] may provide a principled way to select the most informative experiments sequentially.
@@
5 Acknowledgements
We acknowledge Dr Hiroshi Hamamatsu from Sumitomo Chemicals Corporation for producing the OLED device samples whilst on secondment to the University of Sheffield. We are also indebted to Cambridge Display Technologies (CDT) for synthesising and supplying the deuterated conjugated polymers used in the OLED devices. The OLED data was collected on the ISIS neutron reflectometer instrument SURF, and we are grateful to ISIS for beamtime at ISIS under beamtime application RB 1120359. SB, PF and AJP were supported via an EPSRC prosperity partnership, called Sustainable Coatings by Rational Design, SusCoRD, EP/S004963/1, this was jointly funded by the company AkzoNobel. We are grateful for their support and encouragement of this work. This work was (partially) supported by the Centre for Machine Intelligence at the University of Sheffield.
@@ Appendix A Posterior distributions
Posterior distributions identified in this study are included here. In order to limit the size of the figures, only the distributions of the slab model parameters are included in the corner plots.
@@
Figure 10: Identified SVI posterior distributions of NR parameters for the LED data. (a) Background level (shared between all devices). (b) Smearing parameter (shared between all devices). (c) Scale parameter.
@@ Appendix B Benchmark performance on a lipid bilayer system
This section presents results on a benchmark bilayer lipid system that has been previously studied in other published software for the analysis of NR data [9]. As such, this system makes an ideal benchmark for comparison of new fitting methodologies.
The data comprises three NR curves from a lipid bilayer system with different solvents. The solvents considered are D2O, H2O and mixed solvent of D20 and H20. Figure 11 depicts the slab model that is used both in [9] and here to represent the system.
The slab model parameters are derived from the properties of the inner and outer lipid leaflets, the silicon backing, the oxide layer and the solvent. The details of this parameterisation are identical to those in [9] and so are omitted here. Overall there are 78 parameters in the model, however many of these are assumed to be known in the analysis. The authors note that in the approach presented here, all three solvent SLDs are treated as unknown and a different background and scale parameter are inferred for each solvent. In total this approach yields 17 free parameters (6 corresponding to instrument settings and 11 corresponding to the slab model) rather than the 13 free parameters considered in [9].
@@
As can be seen in the figures the fit to the data is very strong in both cases ($\chi^{2}<3$ in all cases). It can also be seen in the results that the posterior variance is slightly lower in the VI result. This is not unexpected and is a known tendency of VI approaches to be ‘mode-seeking’.
@@ Appendix C Remark on the underprediction of variance in VI
Although this variance underprediction is inconvenient, it is sometimes worth the massive computational efficiency savings to use VI over a sampling-based scheme. Ultimately, this decision will come down to how the propagated uncertainty is to be used. If one is only intending to propagate mean and variances of a few parameters through to further analyses (for example fitting a model to a changing parameter in a kinetic run) then VI may well be sufficient. However, for full characterisation of unknown systems it may be more appropriate to sure a ’full-fat’ sampling approach.

### v2

@@ 1.6 Related work
In several works, neural-network surrogates have been proposed as an alternative to optimisation for the NR inverse problem in order to extract unknown parameters directly from data. Methods have been proposed in both NR [doucet2021machine] and X-ray diffraction [pithan2023closing], with the latter motivated as a way to provide real-time process control based on surrogate inversions. A recent review of ML technology applied to scattering problems can be found in [hinderhofer2023machine]. In such cases the principal idea is to pre-train a neural-network surrogate on indicative reflectometry curves (either simulated or experimental) such that it is able to directly extract the parameters of interest. The key advantage is that for a new unseen testing reflectometry curve, the parameters can be extracted in ‘one-shot’ without the need for potentially computationally expensive algorithms when solving the NR inverse problem. While such approaches have merit, it is the opinion of the authors that they are weakened by the requirement to train the methods on large quantities of appropriate labelled training data (experimental or otherwise). While there are certainly situations in which this will be plausible it represents a major limitation to the practicality of such methods.
Recently, surrogate modelling has been proposed in an inference setting for NR. In [starostin2025fast], Starostin et al. propose to apply techniques from neural posterior estimation and prior amortisation to train a neural network to produce credible prior distributions based on reflectivity curves. The idea is that the network is trained to rapidly identify credible regions of the parameter space in the form of prior distributions that can be supplied to traditional Bayesian inference techniques such as Markov-chain Monte-Carlo (MCMC) or importance sampling (IS). The result is a significant reduction in the cost of evaluating a posterior over the NR parameters. The results of the paper are impressive and show that credible prior distributions can be found that include the multimodality that occurs due to the phase invariance problem when using very wide priors. As with all simulation based modelling approaches however, the requirement to pre-train a model (in this case on simulated data from a fixed number of layers) remains a limitation of the approach.
@@
Rather than implement any single bespoke gradient-based optimisation or inference technique for NR, the authors present here an open-source software library for fast computation of forward reflectivity and gradients in the python language. The library, titled refjax, is available at github.com/MDCHAMP/refjax. The library is built on top of the python JAX ecosystem from Google [jax2018github] that provides flexible automatic differentiation, parallelism and GPU acceleration. refjax provides simple access to the forward specular reflectivity calculation (including smearing) for an arbitrary slab model. In refjax, all input parameters are collected in standard python dictionaries with a standardised format.
@@
A key advantage of the provided library is the freedom to apply any optimisation routine developed in the JAX ecosystem and beyond. To date, highly performant packages are available for traditional optimisers (LM, BFGS) [rader2024optimistix], (stochastic) gradient descent [deepmind2020jax], HMC (including the no U-turn sampler [hoffman2014no]) [phan2019composable] and others. Several examples of how to use the library in conjunction with popular ML optimisation suites are provided at https://github.com/MDCHAMP/refjax/tree/main/examples.
@@ 3.1 Crystalline quartz
In the first case study, the performance of gradient based techniques are highlighted on an existing dataset. To this end, the parameters of a highly homogeneous crystal quartz film deposited on a flat silicon wafer with a thin native oxide layer are considered. The reflectivity data were collected on the D17 neutron reflectivity instrument at the ILL (Grenoble, France) [gutfreund2018towards]. As part of the data collection, pointwise full-width at half maximum (FWHM) smearing kernel estimates were estimated. These were used to perform smearing in the forward reflectivity model as fixed known parameters. The slab model considered in this case study is provided in Table 1 and the NR parameters are displayed in Table 2. All parameters are considered unknown and are fit using both the optimisation and inference methods.
@@
First, an optimisation approach is considered. To perform the model fit, the widely-used stochastic gradient descent technique ADAM [kingma2014adam] is employed. Starting from an approximate initial values for the relevant parameters (listed in Tables 1 and 2), the optimiser descends the gradient of a $\chi^{2}$ error metric,
@@
for 2000 iterations with a learning rate of 0.05. The resulting fit to the data is given in Figure 3. As can be seen in the Figure, the fit to the data is excellent, achieving a $\chi^{2}$ error of 1.302. This is marginally lower than the value of $\chi^{2}=1.32$ reported in [gutfreund2018towards], indicating that the approach has found optimal parameters. Note that the best fit parameters for the native silicon thickness, SLD and the substrate SLD do not correspond to physical values in this case, as the adjacent roughnesses are very large. These are just fitting parameters to produce the final best fit SLD profile using a slab model.
Figure 3: Optimisation of quartz thin film NR dataset fit using gradient decent with the ADAM optimiser [kingma2014adam]. All parameters are floating and so free to fit. The identified $\chi^{2}$ error is lower than the value of $\chi^{2}=1.32$ reported previously in [gutfreund2018towards] for the same data.
With the effectiveness of gradient based optimisation established, a further study considering gradient-powered inference for NR fitting is presented here. The same quartz data is used to perform Bayesian inference of the unknown parameters by HMC. The NUTS sampler [hoffman2014no] (as implemented in the ‘numpyro’ library [phan2019composable]) is used to sample from the posterior distribution over the parameters. NUTS has the advantage that the hyperparameter $L$ is set automatically in each iteration leading to efficient exploration of the typical set. The other HMC parameters are adapted using the default strategy in the numypro library for a burn-in period of 1000 samples.
The prior distributions were set to be uniform in the vicinity of the initial parameters used by the gradient scheme (see Tables 1 and 2). The chain was initialised prior to the burn-in at the optimal value found by gradient descent. Overall, 10 independent chains were run, each comprising 2000 samples drawn after burn-in. Posterior means and standard deviations are recorded in Tables 1 and 2. The resulting predictive posterior $p(\hat{R}|R)$ is plotted in Figure 4, wherein an excellent fit to the data is seen. The mean prediction of the posterior $E[\hat{R}$] has a $\chi^{2}$ error of 1.22, an improvement over previously published results ($\chi^{2}=1.32$, [gutfreund2018towards]) although we note that in our work, all parameters are left free to fit. In Figure 5, the marginal posterior distributions of the NR parameters of interest are shown alongside the distribution over the SLD profile of the quartz film.
@@
It is noteworthy that the obtained results are obtained by drawing only 2000 samples from the NUTS algorithm per chain, a fraction of the number of samples typically required by MCMC methods for good convergence. This phenomenon is in line with the efficiency gains expected from HMC over traditional MCMC methods, as demonstrated in other fields [hajian2007efficient, tinto2018gravitational]. However, it is of interest to quantify this phenomenon in the context of NR inversion.
In order to compare the efficiency of HMC (here NUTS) against gradient-free MCMC methods, Figure 6 plots the effective sample size (ESS) against iteration for both HMC and a Sample-adaptive MCMC (SA-MCMC) approach [zhu2019sample] implemented in the numpyro library [phan2019composable]. ESS gives a measure of how many independent samples a correlated Markov-chain is equivalent to. A common rule of thumb in the Bayesian inference literature is that an ESS value of 200 per parameter represents a good coverage of the posterior for well-converged chains, whereas low values can correspond to poor exploration and autocorrelation.
@@
Figure 6: Comparison of effective sample sizes per iteration between a Hamiltonian Monte-Carlo (NUTS) scheme (green) and a gradient-free sample adaptive MCMC approach (orange). Each line is the effective sample size of a single parameter in a single chain. The horizontal line corresponds to a sample size of 200 which is a common rule of thumb for well converged chains in the literature [lanfear2016estimating].
As can be seen in Figure 6, the NUTS samples increase in ESS broadly monotonically with iteration, indicating good convergence and efficient exploration of the typical set even after only 1000 samples. SA-MCMC appears to start well (by exploring a local region of the typical set efficiently) but slow mixing and high auto-correlation lead to a decrease in ESS. It is interesting to note that although the per-sample computational complexity for NUTS is several times higher than SA-MCMC, better exploration of the typical set is seen with the gradient-based scheme, in line with results from other fields [hajian2007efficient, tinto2018gravitational].
@@ 3.2 Organic LED
To produce an approximate inference over the co-fit posterior of all four devices a VI scheme is adopted. The surrogate posterior $q$ is selected to be a multivariate normal distribution defined in an unconstrained space. That is, all parameters are mapped from their physical bounded space to one that is unconstrained by an appropriate invertible transformation. The resulting ELBO is evaluated by automatic differentiation (using the numpyro library) and the refjax software for reflectivity kernel calculations. The mean of the surrogate posterior is initialised at identical values for each device and the variance is initialised as a diagonal matrix $\sigma_{0}I$ with $\sigma_{0}=0.1$. The ELBO is optimised using the ADAM [kingma2014adam] optimiser for 2000 iterations with a learning rate of 0.05. To ensure that the optimisation does not become trapped in local minima, the optimisation procedure is run 100 times and the best (largest) value of the ELBO is retained as surrogate posterior. VI has significant computational advantages over sampling based approaches, each run of the inference is complete in <20 seconds on our hardware (AMD Ryzen Threadripper PRO 5995WX). Successive restarts can be run entirely in parallel and so the VI approach is very attractive with regard to modern CPU and GPU acceleration.
@@ 4 Discussion
In the second case-study, techniques from variational inference show remarkable performance in a complex NR fitting task. The authors believe that VI can be revolutionary for several NR applications particularly in settings whereby uncertainty quantification of high-throughput data is critical. An example is fast kinetic NR experiments whereby large quantities of NR data are collected dynamically and in-situ inference of model parameters is required. A key application of VI in NR might be in real-time parameter inference for closed-loop process control, something that has thus far required fast model surrogates and has been unattainable with traditional inverse modelling techniques. Another application could be in iterative experimental design whereby the parameter posterior distributions of initial experiments are used to inform the design of subsequent experiments during a single beamtime. Here, techniques from Bayesian optimisation [ru2018fast] may provide a principled way to select the most informative experiments sequentially.
@@
Author contributions
MDC: Conceptualization, Data curation, Writing - Original Draft, Visualization, Software; AJP: Conceptualization, Data curation, Writing - Review & Editing, Funding acquisition; PG: Data curation, Writing - Review & Editing; MAS: Writing - Review & Editing; JPAF: Funding acquisition, Writing - Review & Editing; TJR: Conceptualization, Writing - Review & Editing, Funding acquisition; SLB: Conceptualization, Funding acquisition, Writing - Review & Editing;
Funding statement
SB, PF and AJP were supported via an EPSRC prosperity partnership, called Sustainable Coatings by Rational Design, SusCoRD, EP/S004963/1, this was jointly funded by the company AkzoNobel.
The OLED data was collected on the ISIS neutron reflectometer instrument SURF, and we are grateful to ISIS for beamtime at ISIS under beamtime application RB 1120359. We are grateful for their support and encouragement of this work.
Competing interests
We confirm that there are no competing interests to declare.
Acknowledgements
We acknowledge Dr Hiroshi Hamamatsu from Sumitomo Chemicals Corporation for producing the OLED device samples whilst on secondment to the University of Sheffield. We are also indebted to Cambridge Display Technologies (CDT) for synthesising and supplying the deuterated conjugated polymers used in the OLED devices. This work was (partially) supported by the Centre for Machine Intelligence at the University of Sheffield.
@@ Appendix A Posterior distributions
Posterior distributions identified in this study are included here. In order to limit the size of the figures, only the distributions of the slab model parameters are included in the corner plots. Posterior distributions for the quartz and LED data are plotted in Figures 8 and 9 respectively. Distributions of the instrument parameters are plotted in Figure 10 Posterior parameter values from the LED data are collected in 5 for the slab model parameters and in 6 for the instrument parameters.
@@
Figure 10: Identified VI posterior distributions of NR parameters for the LED data. (a) Background level (shared between all devices). (b) Smearing parameter (shared between all devices). (c) Scale parameter.
@@ Appendix B Benchmark performance on a lipid bilayer system
This section presents results on a benchmark bilayer lipid system that has been previously studied in other published software for the analysis of NR data [nelson2019refnx]. As such, this system makes an ideal benchmark for comparison of new fitting methodologies.
The data comprises three NR curves from a lipid bilayer system with different solvents. The solvents considered are D2O, H2O and mixed solvent of D20 and H20. Figure 11 depicts the slab model that is used both in [nelson2019refnx] and here to represent the system.
The slab model parameters are derived from the properties of the inner and outer lipid leaflets, the silicon backing, the oxide layer and the solvent. The details of this parameterisation are identical to those in [nelson2019refnx] and so are omitted here. Overall there are 78 parameters in the model, however many of these are assumed to be known in the analysis. The authors note that in the approach presented here, all three solvent SLDs are treated as unknown and a different background and scale parameter are inferred for each solvent. In total this approach yields 17 free parameters (6 corresponding to instrument settings and 11 corresponding to the slab model) rather than the 13 free parameters considered in [nelson2019refnx].
@@
As can be seen in Figure 12 the fit to the data is very strong in both cases ($\chi^{2}<3$ in all cases). It can also be seen in Figures 13 and 14 that the posterior variance is slightly lower in the VI result. This is not unexpected and is a known tendency of VI approaches to be ‘mode-seeking’.
@@ Appendix C Remark on the underprediction of variance in VI
Although this variance underprediction is inconvenient, it is sometimes worth the massive computational efficiency savings to use VI over a sampling-based scheme. Ultimately, this decision will come down to how the propagated uncertainty is to be used. If one is only intending to propagate mean and variances of a few parameters through to further analyses (for example fitting a model to a changing parameter in a kinetic run) then VI may well be sufficient. However, for full characterisation of unknown systems it may be more appropriate to use a ’full-fat’ sampling approach.